Solve the initial value problem y"-3y'-2y=0; y(0)=1, y(3)=0


1
Expert's answer
2021-12-07T10:02:41-0500

The auxiliary equation is "m^2-3m-2=0"

"m= \dfrac{3\pm \sqrt{9-4(1)(-2)}}{2} = \dfrac{3 \pm \sqrt{17}}{2}"

"m = \dfrac{3+\sqrt{17}}{2}" Or "m = \dfrac{3-\sqrt{17}}{2}"

"m = 3.56" or "m = -0.56"

Hence the general solution is "y = Ae^{3.56x} + Be^{-0.56x}"

So "1= y(0) = A+B .... (*)"

And "0= y(3)= 43477.55A+0.19B .... (**)"

So "B = \dfrac{-43477.55A}{0.19} = - 228829.21A"

From equation (*) 1=A−228829.21A=−228828.21A

So A =−4.37×10"^{-6}"

And B = 0.99998

So the solution of the IVP is y = "-4.37 \times 10^{-6}e^{3.56x} + 0.99998e^{-0.56x}"



NOTE: Although what I have written down is the logical solution for the prescribed question. But I think there's a mistake in the question especially with respect to the boundary condition.